# Computing Balanced Solutions for Large International Kidney Exchange Schemes When Cycle Length Is Unbounded

### v1

@@ 5 Conclusions
We introduced the class of partitioned permutation games and proved a number of complexity results that contrast known results for partitioned matching games. Our new results guided our simulations for IKEPs with up to ten countries, with exchange bound $\ell=\infty$. Our simulations showed a significant improvement in the total number of kidney transplants over the case where $\ell=2$ [8] at the expense of long cycles. In our simulations, all countries had the same size. In [8], simulations were also done for countries with three different sizes, but these led to the same conclusions. We expect the same for $\ell=\infty$, as confirmed by a robustness check for $n=6$ (see Github repository [42]).
@@
For future research we will consider the more realistic exchange bounds $\ell\in\{3,4,5\}$. We note that the simulations done in [11, 31] were only for $\ell=3$; a more limited number of solution concepts; for IKEPs with up to four countries; and only for scenarios that use weakly close optimal solutions. They also used different data sets. For our follow-up study for $\ell\in\{3,4,5\}$, we must now also overcome, just like [11, 31], the additional computational obstacle of not being able to compute an optimal solution for a compatibility graph in a KEP round and the values $v(S)$ of the associated permutation game in polynomial time (see Theorem 1.1). Current techniques for $\ell\in\{3,4,5\}$ therefore involve, besides ILPS based on the edge-formulation, ILPs based on the cycle-formulation, with a variable for each cycle of length at most $\ell$ (see, for example, [20, 22, 25]). Hence, computing a single value $v(S)$ will become significantly more expensive, and even more so for increasing $\ell$. For expensive solution concepts, such as the Shapley or nucleolus, we must compute an exponential number of values $v(S)$. Without any new methods, we expect it will not be possible to do this for simulations up to the same number of countries (ten) as we did for $\ell=\infty$.
@@
Finally, we note that matching games and permutation games are usually defined on edge-weighted graphs. It is readily seen that all our positive theoretical results can be generalized to this setting. In kidney exchange, the primary goal is still to help as many patients as possible, but edge weights might be used to represent transplant utilities. Hence, it would also be interesting to do simulations in the presence of edge weights. We leave this as future research.
Acknowledgments. Benedek was supported by the National Research, Development and Innovation Office of Hungary (OTKA Grant No. K138945); Biró by the Hungarian Scientific Research Fund (OTKA, Grant No. K143858) and the Hungarian Academy of Sciences (Momentum Grant No. LP2021-2); Csáji by the Hungarian Scientific Research Fund (OTKA Grant No. K143858) and the Momentum Grant of the Hungarian Academy of Sciences (Grant No. 2021-1/2021); and Paulusma was supported by the Leverhulme Trust (Grant RF-2022-607) and EPSRC (Grant EP/X01357X/1). Moreover, this work has used Durham University’s NCC cluster. NCC has been purchased through Durham University’s strategic investment funds, and is installed and maintained by the Department of Computer Science. In particular, we thank Rob Powell for his help with setting up our simulations on the NCC cluster.

### v2

@@ 5 Conclusions
We introduced the class of partitioned permutation games and proved a number of complexity results that contrast known results for partitioned matching games. As open theoretical problems we ask: what is the complexity of computing the nucleolus for permutation games and for partitioned permutation games?
@@
Our new results guided our simulations for IKEPs with up to ten countries, with exchange bound $\ell=\infty$. Our simulations showed a significant improvement in the total number of kidney transplants over the case where $\ell=2$ [10] at the expense of long cycles. In our simulations, we first let all countries be of the same size. We then examined in the same way as done in [10], whether our conclusions would change for varying country sizes. As in [10], this turned out not to be the case.
@@
For future research we will consider the more realistic exchange bounds $\ell\in\{3,4,5\}$. We note that the simulations done in [12, 32] were only for $\ell=3$; a more limited number of solution concepts; for IKEPs with up to four countries; and only for scenarios that use weakly close optimal solutions. They also used different data sets. For our follow-up study for $\ell\in\{3,4,5\}$, we must now also overcome, as in [12, 32], the additional computational obstacle of not being able to compute an optimal solution for a compatibility graph in a KEP round and the values $v(S)$ of the associated permutation game in polynomial time (see Theorem 1.1). Current techniques for $\ell\in\{3,4,5\}$ therefore involve, besides ILPs based on the edge-formulation, ILPs based on the cycle-formulation, with a variable for each cycle of length at most $\ell$ (see, for example, [21, 23, 26]). Hence, computing a single value $v(S)$ will become significantly more expensive, and even more so for increasing $\ell$. For expensive solution concepts, such as the Shapley or nucleolus, we must compute an exponential number of values $v(S)$. Without any new methods, we expect it will not be possible to do this for simulations up to the same number of countries (ten) as we did for $\ell=\infty$.
@@
Finally, we note that matching games and permutation games are usually defined on edge-weighted graphs. It is readily seen that all our positive theoretical results can be generalized to this setting. In kidney exchange, the primary goal is still to help as many patients as possible, but edge weights might be used to represent transplant utilities. Hence, it would also be interesting to do simulations in the presence of edge weights. We leave this for future research as well.
Acknowledgments. Benedek was supported by the National Research, Development and Innovation Office of Hungary (OTKA Grant No. K138945); Biró by the Hungarian Scientific Research Fund (OTKA, Grant No. K143858) and the Hungarian Academy of Sciences (Momentum Grant No. LP2021-2); Csáji by the Hungarian Scientific Research Fund, OTKA, Grant No. K143858, the ÚNKP-23-3 New National Excellence Program of the Ministry for Culture and Innovation and by the Doctoral Student Scholarship Program (number C2258525) of the Cooperative Doctoral Program of the Ministry of Innovation and Technology financed by the National Research, Development, and Innovation Fund; and Paulusma was supported by the Leverhulme Trust (Grant RF-2022-607) and EPSRC (Grant EP/X01357X/1).
This work has made use of the Hamilton HPC Service of Durham University. In particular, we thank Rob Powell for all his help with the initial set up of our simulations.
Appendix 0.A Average Maximum Relative Deviations
In this section we consider the average maximum relative deviation as our evaluation measure. We refer to Figures 1215 for similar figures as in Section 4 where we used the average total relative deviation.
Figure 12: Average maximum relative deviations for each of the seven solution concepts under the five different scenarios for equal country sizes, where the number of countries $n$ is ranging from $4$ to $10$. For comparison, the lower right figure displays a result from [10] for $\ell=2$, namely for the Banzhaf* value, which behaved best for $\ell=2$.
Figure 13: Average maximum relative deviations for each of the seven solution concepts under the five different scenarios for varying country sizes, where the number of countries $n$ is ranging from $4$ to $10$. For comparison, the lower right figure displays a result from [10] for $\ell=2$, namely for the Banzhaf* value, which behaved best for $\ell=2$.
Figure 14: Average maximum relative deviations for all solution concepts in the lexmin+c scenario, where the number of countries $n$ ranges from $4$ to $10$.
Figure 15: Average maximum relative deviations for all solution concepts in the d1+c scenario, where the number of countries $n$ ranges from $4$ to $10$.
Appendix 0.B Additional Information on Incomplete Simulations
#unsolved / n 4 5 6 7 8 9 10 Total
Shapley value
d1 0 0 0 0 0 0 0 0
d1+c 0 0 0 0 0 0 0 0
lexmin 0 0 0 3 2 14 7 26
lexmin+c 0 0 0 3 5 13 7 28
Nucleolus
d1 0 0 0 0 0 0 0 0
d1+c 0 0 0 0 0 0 0 0
lexmin 0 0 0 0 0 0 0 0
lexmin+c 0 0 0 2 0 0 1 3
Benefit value
d1 0 0 0 0 0 0 0 0
d1+c 0 0 0 0 0 0 0 0
lexmin 0 0 0 0 1 0 0 1
lexmin+c 1 1 5 1 2 3 0 13
Contribution value
d1 0 0 0 0 0 0 0 0
d1+c 0 0 0 0 0 0 0 0
lexmin 1 1 2 1 1 2 4 12
lexmin+c 2 2 5 3 5 0 5 22
Tau value
d1 0 0 0 0 0 1 0 1
d1+c 0 0 0 0 0 1 0 1
lexmin 0 0 0 0 0 1 0 1
lexmin+c 0 1 0 0 0 1 0 2
Banzhaf value
d1 0 0 0 0 0 0 0 0
d1+c 0 0 0 0 0 0 0 0
lexmin 0 1 1 4 1 8 9 24
lexmin+c 0 2 5 3 1 9 9 29
Banzhaf* value
d1+c 0 0 0 0 0 0 0 0
lexmin+c 0 0 0 0 0 0 1 1
Table 5: Complete breakdown of the number of incomplete simulation instances (out of 100) for the seven different solution concepts and the four different scenarios (excluding the arbitrary matching scenario, where no ILPs were used) for equal country sizes.
#unfinished / n 4 5 6 7 8 9 10 Total
Shapley value
d1 0 0 0 0 0 0 0 0
d1+c 0 0 0 0 0 0 0 0
lexmin 0 0 0 1 0 0 6 7
lexmin+c 0 0 0 1 0 2 2 5
Nucleolus
d1 0 0 0 0 0 1 0 1
d1+c 0 0 0 0 0 0 0 0
lexmin 0 0 0 0 0 1 0 1
lexmin+c 0 0 0 0 0 1 0 1
Benefit value
d1 0 0 0 0 0 0 0 0
d1+c 0 0 1 0 0 0 0 1
lexmin 0 0 0 0 0 0 0 0
lexmin+c 1 0 2 0 0 1 3 7
Contribution value
d1 0 0 0 0 0 0 0 0
d1+c 0 0 0 0 0 0 0 0
lexmin 0 0 1 0 1 1 0 3
lexmin+c 0 0 1 0 1 1 1 4
Tau value
d1 0 0 0 0 0 0 0 0
d1+c 0 0 0 0 1 0 0 1
lexmin 0 0 0 0 0 0 0 0
lexmin+c 0 0 0 0 0 0 0 0
Banzhaf value
d1 0 1 0 0 0 0 0 1
d1+c 0 1 0 0 0 0 0 1
lexmin 0 1 1 1 2 1 4 10
lexmin+c 0 1 0 0 3 1 4 9
Banzhaf* value
d1+c 0 1 0 0 0 0 0 1
lexmin+c 0 1 0 0 1 1 4 7
Table 6: Complete breakdown of the number of incomplete simulation instances (out of 100) for the seven different solution concepts and the four different scenarios (excluding the arbitrary matching scenario, where no ILPs were used) for varying country sizes.
Figure 16: Average total relative deviations, leaving out the incomplete simulation instances under the five different scenarios for equal country sizes, where the number of countries ranges from $4$ to $10$. We note that this figure is identical to Figure 4 that includes all simulation instances.
Figure 17: Average total relative deviations, leaving out the incomplete simulation instances under the five different scenarios for varying country sizes, where the number of countries ranges from $4$ to $10$. We note that this figure is identical to Figure 5 that includes all simulation instances.
Appendix 0.C Relative Ratios
n 4 5 6 7 8 9 10
Shapley value $\ell=\infty$ 44.43% 40.58% 37.71% 33.29% 32.18% 31.99% 31.07%
$\ell=2$ 42.82% 38.62% 35.31% 31.05% 27.83% 26.96% 24.81%
difference 1.61% 1.96% 2.40% 2.24% 4.35% 5.03% 6.26%
Banzhaf value $\ell=\infty$ 43.75% 39.53% 40.34% 35.92% 32.83% 31.31% 30.06%
$\ell=2$ 43.18% 37.15% 33.94% 31.19% 30.13% 27.61% 24.50%
difference 0.57% 2.38% 6.40% 4.73% 2.70% 3.70% 5.56%
nucleolus $\ell=\infty$ 43.99% 37.67% 37.67% 35.10% 29.67% 30.08% 29.06%
$\ell=2$ 46.06% 39.05% 35.94% 34.13% 31.35% 30.26% 27.10%
difference -2.07% -1.38% 1.73% 0.98% -1.68% -0.18% 1.96%
tau value $\ell=\infty$ 45.63% 42.38% 39.61% 34.47% 31.25% 30.81% 28.82%
$\ell=2$ 41.92% 38.50% 35.39% 36.04% 31.29% 28.05% 26.06%
difference 3.72% 3.88% 4.22% -1.57% -0.04% 2.77% 2.76%
benefit value $\ell=\infty$ 45.53% 40.96% 40.33% 34.19% 32.73% 29.63% 27.95%
$\ell=2$ 42.31% 37.14% 33.68% 31.91% 30.41% 28.76% 25.05%
difference 3.22% 3.82% 6.65% 2.28% 2.33% 0.87% 2.91%
contribution value $\ell=\infty$ 39.67% 36.83% 36.54% 35.61% 31.01% 29.51% 27.45%
$\ell=2$ 43.57% 36.74% 36.41% 32.65% 28.38% 27.49% 25.31%
difference -3.91% 0.09% 0.13% 2.96% 2.63% 2.02% 2.15%
Banzhaf* value $\ell=\infty$ 44.26% 40.10% 38.69% 36.31% 31.90% 31.17% 29.39%
$\ell=2$ 41.60% 35.21% 31.94% 28.43% 27.12% 22.61% 21.53%
difference 2.66% 4.89% 6.75% 7.88% 4.78% 8.56% 7.86%
Table 7: Relative ratios for varying country sizes under lexmin+c, for $\ell=\infty$, $\ell=2$ and their difference.
n 4 5 6 7 8 9 10
Shapley value $\ell=\infty$ 39.35% 30.99% 29.02% 25.07% 22.48% 21.05% 18.76%
$\ell=2$ 43.38% 38.19% 33.14% 30.28% 26.48% 23.92% 22.07%
difference -4.03% -7.19% -4.12% -5.21% -4.00% -2.87% -3.31%
Banzhaf value $\ell=\infty$ 39.74% 33.02% 29.33% 24.52% 22.23% 20.96% 18.82%
$\ell=2$ 42.73% 35.31% 30.25% 27.88% 25.00% 22.51% 21.07%
difference -2.99% -2.29% -0.92% -3.36% -2.77% -1.55% -2.25%
nucleolus $\ell=\infty$ 40.64% 32.40% 29.62% 25.55% 23.57% 21.75% 19.80%
$\ell=2$ 42.33% 37.23% 31.97% 29.21% 25.00% 23.09% 20.56%
difference -1.69% -4.83% -2.36% -3.65% -1.43% -1.33% -0.76%
tau value $\ell=\infty$ 39.97% 32.84% 29.41% 24.10% 23.24% 21.48% 19.39%
$\ell=2$ 56.85% 46.43% 37.42% 31.95% 29.29% 26.97% 24.74%
difference -16.88% -13.59% -8.01% -7.85% -6.05% -5.49% -5.35%
benefit value $\ell=\infty$ 39.59% 32.02% 29.45% 24.35% 23.49% 21.67% 19.32%
$\ell=2$ 41.36% 38.72% 33.50% 28.94% 26.86% 23.02% 20.49%
difference -1.77% -6.71% -4.05% -4.59% -3.37% -1.35% -1.18%
contribution value $\ell=\infty$ 39.79% 31.56% 29.06% 25.00% 22.37% 20.50% 18.45%
$\ell=2$ 42.38% 38.24% 33.41% 29.89% 26.22% 23.84% 21.14%
difference -2.59% -6.68% -4.36% -4.89% -3.85% -3.33% -2.69%
Banzhaf* value $\ell=\infty$ 41.21% 34.16% 28.94% 26.12% 23.40% 20.99% 19.57%
$\ell=2$ 42.40% 35.79% 31.21% 28.45% 24.70% 22.88% 21.08%
difference -1.19% -1.62% -2.27% -2.33% -1.30% -1.89% -1.51%
Table 8: Relative ratios for equal country sizes under d1+c, for $\ell=\infty$, $\ell=2$ and their difference.
n 4 5 6 7 8 9 10
Shapley value $\ell=\infty$ 39.18% 31.66% 27.75% 25.12% 22.51% 21.39% 18.83%
$\ell=2$ 42.00% 37.21% 31.11% 27.88% 25.48% 22.45% 19.78%
difference -2.82% -5.56% -3.35% -2.76% -2.98% -1.05% -0.95%
Banzhaf value $\ell=\infty$ 38.21% 33.81% 30.03% 25.03% 22.51% 20.96% 19.55%
$\ell=2$ 41.64% 37.85% 30.53% 27.92% 25.37% 22.68% 21.09%
difference -3.42% -4.04% -0.49% -2.89% -2.85% -1.72% -1.54%
nucleolus $\ell=\infty$ 34.29% 28.92% 26.69% 25.16% 21.51% 20.13% 19.42%
$\ell=2$ 39.79% 32.16% 26.26% 23.07% 21.66% 20.33% 18.24%
difference -5.50% -3.24% 0.43% 2.10% -0.15% -0.20% 1.17%
tau value $\ell=\infty$ 37.66% 32.07% 29.60% 25.25% 22.80% 21.94% 19.10%
$\ell=2$ 46.67% 35.53% 30.91% 29.27% 24.75% 21.81% 20.12%
difference -9.01% -3.45% -1.31% -4.02% -1.96% 0.13% -1.02%
benefit value $\ell=\infty$ 37.56% 33.48% 29.28% 25.08% 22.56% 21.38% 19.45%
$\ell=2$ 41.50% 34.41% 32.13% 26.41% 24.49% 21.74% 19.71%
difference -3.94% -0.93% -2.85% -1.32% -1.93% -0.35% -0.26%
contribution value $\ell=\infty$ 31.23% 29.10% 26.19% 25.19% 21.45% 19.59% 18.38%
$\ell=2$ 38.51% 32.26% 28.71% 25.60% 22.88% 20.89% 19.32%
difference -7.27% -3.16% -2.53% -0.42% -1.42% -1.30% -0.94%
Banzhaf* value $\ell=\infty$ 38.21% 33.81% 30.03% 25.03% 22.51% 20.96% 19.55%
$\ell=2$ 41.64% 37.85% 30.53% 27.92% 25.37% 22.68% 21.09%
difference -3.42% -4.04% -0.49% -2.89% -2.85% -1.72% -1.54%
Table 9: Relative ratios for varying country sizes under d1+c, for $\ell=\infty$, $\ell=2$ and their difference.
n 4 5 6 7 8 9 10
Shapley value $\ell=\infty$ 43.66% 36.50% 35.01% 31.87% 27.73% 25.86% 23.59%
$\ell=2$ 43.02% 37.65% 34.21% 30.58% 27.36% 25.96% 23.11%
difference 0.64% -1.15% 0.80% 1.29% 0.37% -0.10% 0.48%
Banzhaf value $\ell=\infty$ 43.04% 36.34% 35.15% 31.14% 27.69% 25.74% 22.87%
$\ell=2$ 43.71% 36.90% 33.85% 29.67% 28.06% 25.64% 23.42%
difference -0.66% -0.56% 1.31% 1.48% -0.37% 0.10% -0.56%
nucleolus $\ell=\infty$ 43.97% 36.83% 34.53% 31.64% 28.27% 25.31% 23.27%
$\ell=2$ 44.57% 36.52% 34.48% 29.52% 27.85% 26.20% 24.17%
difference -0.59% 0.31% 0.06% 2.12% 0.41% -0.89% -0.90%
tau value $\ell=\infty$ 43.55% 37.36% 34.61% 31.16% 27.79% 24.95% 23.68%
$\ell=2$ 44.00% 39.67% 33.87% 30.85% 26.95% 25.18% 23.49%
difference -0.45% -2.31% 0.73% 0.31% 0.84% -0.23% 0.18%
benefit value $\ell=\infty$ 42.98% 37.59% 33.34% 32.33% 28.17% 24.73% 24.22%
$\ell=2$ 42.98% 39.41% 32.90% 30.28% 26.75% 25.73% 23.74%
difference 0.00% -1.83% 0.44% 2.05% 1.42% -1.00% 0.47%
contribution value $\ell=\infty$ 42.92% 35.82% 34.11% 30.26% 27.68% 24.79% 23.13%
$\ell=2$ 44.52% 37.22% 33.77% 31.10% 28.12% 26.39% 22.57%
difference -1.60% -1.40% 0.33% -0.84% -0.44% -1.60% 0.56%
Banzhaf* value $\ell=\infty$ 43.04% 36.34% 35.15% 31.14% 27.69% 25.74% 22.87%
$\ell=2$ 43.71% 36.90% 33.85% 29.67% 28.06% 25.64% 23.42%
difference -0.66% -0.56% 1.31% 1.48% -0.37% 0.10% -0.56%
Table 10: Relative ratios for equal country sizes under lexmin, for $\ell=\infty$, $\ell=2$ and their difference.
n 4 5 6 7 8 9 10
Shapley value $\ell=\infty$ 42.41% 39.85% 35.90% 30.36% 28.87% 25.47% 22.75%
$\ell=2$ 45.17% 37.62% 33.77% 29.46% 27.20% 25.13% 23.66%
difference -2.76% 2.23% 2.14% 0.90% 1.67% 0.34% -0.90%
Banzhaf value $\ell=\infty$ 42.44% 39.50% 34.98% 30.66% 29.63% 25.36% 23.76%
$\ell=2$ 42.08% 39.60% 34.42% 30.31% 26.87% 25.73% 24.03%
difference 0.36% -0.11% 0.56% 0.35% 2.75% -0.37% -0.26%
nucleolus $\ell=\infty$ 40.74% 35.71% 33.04% 30.84% 27.59% 26.51% 23.65%
$\ell=2$ 42.53% 38.17% 34.08% 29.74% 28.51% 26.64% 25.27%
difference -1.80% -2.46% -1.05% 1.10% -0.92% -0.14% -1.62%
tau value $\ell=\infty$ 41.81% 38.72% 34.16% 30.77% 28.55% 26.31% 23.12%
$\ell=2$ 45.29% 37.62% 34.39% 29.38% 27.49% 25.13% 23.88%
difference -3.48% 1.11% -0.22% 1.38% 1.06% 1.18% -0.76%
benefit value $\ell=\infty$ 42.52% 39.28% 34.44% 30.99% 28.39% 25.26% 23.82%
$\ell=2$ 43.70% 37.10% 33.92% 30.51% 27.40% 25.11% 23.98%
difference -1.18% 2.18% 0.51% 0.47% 0.99% 0.15% -0.16%
contribution value $\ell=\infty$ 36.57% 35.10% 31.46% 31.39% 26.63% 25.40% 23.02%
$\ell=2$ 37.19% 33.40% 32.36% 30.46% 28.79% 27.69% 24.80%
difference -0.63% 1.70% -0.89% 0.93% -2.16% -2.29% -1.78%
Banzhaf* value $\ell=\infty$ 42.44% 39.50% 34.98% 30.66% 29.63% 25.36% 23.76%
$\ell=2$ 42.08% 39.60% 34.42% 30.31% 26.87% 25.73% 24.03%
difference 0.36% -0.11% 0.56% 0.35% 2.75% -0.37% -0.26%
Table 11: Relative ratios for varying country sizes under lexmin, for $\ell=\infty$, $\ell=2$ and their difference.
n 4 5 6 7 8 9 10
Shapley value $\ell=\infty$ 42.66% 37.54% 32.40% 30.34% 26.29% 26.12% 22.79%
$\ell=2$ 42.57% 37.54% 34.20% 30.24% 27.99% 24.22% 22.87%
difference 0.09% 0.00% -1.81% 0.09% -1.70% 1.90% -0.08%
Banzhaf value $\ell=\infty$ 43.36% 37.84% 34.55% 31.11% 26.35% 24.97% 22.70%
$\ell=2$ 43.96% 37.92% 33.42% 29.87% 27.04% 25.00% 23.11%
difference -0.60% -0.08% 1.13% 1.24% -0.69% -0.03% -0.40%
nucleolus $\ell=\infty$ 42.18% 36.51% 33.15% 31.99% 27.25% 24.81% 23.85%
$\ell=2$ 45.00% 37.98% 32.49% 29.57% 27.37% 25.42% 22.73%
difference -2.82% -1.47% 0.66% 2.42% -0.13% -0.61% 1.12%
tau value $\ell=\infty$ 42.54% 36.46% 33.55% 30.46% 27.41% 24.95% 23.71%
$\ell=2$ 42.68% 37.32% 33.21% 29.89% 28.12% 25.07% 23.42%
difference -0.13% -0.86% 0.34% 0.57% -0.71% -0.12% 0.28%
benefit value $\ell=\infty$ 41.25% 37.80% 33.96% 30.01% 27.99% 25.01% 23.54%
$\ell=2$ 43.67% 37.81% 33.68% 29.98% 27.14% 24.64% 23.92%
difference -2.42% -0.01% 0.28% 0.02% 0.85% 0.36% -0.39%
contribution value $\ell=\infty$ 43.32% 36.90% 34.20% 30.21% 27.44% 24.70% 22.31%
$\ell=2$ 42.91% 37.76% 32.70% 30.33% 27.63% 25.55% 23.26%
difference 0.40% -0.86% 1.49% -0.11% -0.19% -0.85% -0.96%
Banzhaf* value $\ell=\infty$ 43.36% 37.84% 34.55% 31.11% 26.35% 24.97% 22.70%
$\ell=2$ 43.96% 37.92% 33.42% 29.87% 27.04% 25.00% 23.11%
difference -0.60% -0.08% 1.13% 1.24% -0.69% -0.03% -0.40%
Table 12: Relative ratios for equal country sizes under d1, for $\ell=\infty$, $\ell=2$ and their difference.
n 4 5 6 7 8 9 10
Shapley value $\ell=\infty$ 42.74% 38.66% 33.65% 31.96% 28.69% 25.47% 23.17%
$\ell=2$ 44.61% 38.56% 33.07% 29.70% 27.57% 24.85% 22.91%
difference -1.87% 0.10% 0.58% 2.26% 1.13% 0.62% 0.26%
Banzhaf value $\ell=\infty$ 44.07% 39.81% 34.52% 31.04% 28.49% 25.85% 23.58%
$\ell=2$ 42.96% 38.70% 34.52% 29.45% 27.90% 25.48% 23.01%
difference 1.11% 1.11% 0.00% 1.60% 0.59% 0.38% 0.57%
nucleolus $\ell=\infty$ 38.99% 35.17% 31.74% 30.03% 26.66% 25.90% 23.07%
$\ell=2$ 41.91% 36.93% 33.97% 29.82% 27.56% 25.70% 23.59%
difference -2.93% -1.75% -2.23% 0.21% -0.89% 0.20% -0.52%
tau value $\ell=\infty$ 41.58% 37.88% 33.09% 31.44% 27.77% 26.33% 23.96%
$\ell=2$ 48.20% 38.31% 32.66% 30.14% 26.17% 25.75% 23.44%
difference -6.62% -0.43% 0.43% 1.30% 1.60% 0.59% 0.52%
benefit value $\ell=\infty$ 42.62% 38.50% 34.37% 31.24% 28.34% 25.48% 24.08%
$\ell=2$ 41.88% 36.39% 34.85% 30.41% 28.05% 25.58% 22.31%
difference 0.75% 2.11% -0.48% 0.84% 0.29% -0.10% 1.76%
contribution value $\ell=\infty$ 36.67% 33.28% 30.37% 30.77% 25.79% 23.44% 21.56%
$\ell=2$ 38.40% 33.79% 31.53% 30.64% 27.91% 25.48% 24.00%
difference -1.73% -0.51% -1.16% 0.13% -2.12% -2.04% -2.44%
Banzhaf* value $\ell=\infty$ 44.07% 39.81% 34.52% 31.04% 28.49% 25.85% 23.58%
$\ell=2$ 42.96% 38.70% 34.52% 29.45% 27.90% 25.48% 23.01%
difference 1.11% 1.11% 0.00% 1.60% 0.59% 0.38% 0.57%
Table 13: Relative ratios for varying country sizes under d1, for $\ell=\infty$, $\ell=2$ and their difference.
n 4 5 6 7 8 9 10
Shapley value $\ell=\infty$ 43.98% 38.25% 33.81% 30.90% 26.32% 25.66% 23.25%
$\ell=2$ 39.36% 34.50% 29.35% 26.18% 23.18% 21.57% 19.49%
difference 4.62% 3.75% 4.46% 4.72% 3.14% 4.09% 3.75%
Banzhaf value $\ell=\infty$ 44.06% 38.36% 33.94% 30.56% 27.16% 25.49% 23.25%
$\ell=2$ 39.37% 34.46% 29.35% 26.26% 23.19% 21.57% 19.55%
difference 4.68% 3.90% 4.59% 4.31% 3.97% 3.92% 3.70%
nucleolus $\ell=\infty$ 43.55% 37.44% 34.60% 30.29% 27.71% 25.34% 24.00%
$\ell=2$ 39.83% 35.14% 29.60% 26.69% 23.49% 22.09% 20.07%
difference 3.71% 2.29% 5.00% 3.59% 4.22% 3.26% 3.93%
tau value $\ell=\infty$ 43.20% 37.75% 34.42% 29.98% 27.80% 25.45% 23.76%
$\ell=2$ 39.50% 34.80% 29.60% 26.25% 23.27% 21.74% 19.72%
difference 3.70% 2.95% 4.82% 3.73% 4.53% 3.71% 4.04%
benefit value $\ell=\infty$ 43.12% 37.85% 34.67% 29.64% 27.81% 24.61% 23.38%
$\ell=2$ 39.39% 34.84% 29.48% 26.04% 23.25% 21.76% 19.59%
difference 3.73% 3.01% 5.19% 3.60% 4.57% 2.85% 3.78%
contribution value $\ell=\infty$ 44.02% 37.40% 34.34% 30.24% 26.69% 24.59% 23.38%
$\ell=2$ 39.56% 34.43% 29.99% 26.29% 23.49% 22.17% 20.12%
difference 4.46% 2.97% 4.35% 3.95% 3.21% 2.43% 3.27%
Banzhaf* value $\ell=\infty$ 44.06% 38.36% 33.94% 30.56% 27.16% 25.49% 23.25%
$\ell=2$ 39.37% 34.46% 29.35% 26.26% 23.19% 21.57% 19.55%
difference 4.68% 3.90% 4.59% 4.31% 3.97% 3.92% 3.70%
Table 14: Relative ratios for equal country sizes under arbitrary, for $\ell=\infty$, $\ell=2$ and their difference.
n 4 5 6 7 8 9 10
Shapley value $\ell=\infty$ 43.61% 39.47% 35.60% 31.60% 29.39% 26.50% 23.95%
$\ell=2$ 48.45% 44.14% 34.83% 33.25% 28.02% 27.37% 23.15%
difference -4.84% -4.67% 0.77% -1.65% 1.38% -0.87% 0.79%
Banzhaf value $\ell=\infty$ 43.90% 39.04% 35.01% 31.64% 29.10% 26.15% 23.91%
$\ell=2$ 47.85% 42.82% 34.13% 32.82% 27.54% 26.71% 22.78%
difference -3.96% -3.78% 0.88% -1.18% 1.56% -0.56% 1.13%
nucleolus $\ell=\infty$ 41.05% 36.85% 33.31% 31.74% 28.16% 26.13% 23.71%
$\ell=2$ 49.57% 47.31% 37.30% 35.07% 29.78% 29.64% 24.81%
difference -8.52% -10.46% -4.00% -3.33% -1.62% -3.50% -1.10%
tau value $\ell=\infty$ 43.03% 38.65% 34.12% 31.94% 29.69% 25.28% 24.72%
$\ell=2$ 54.97% 45.25% 35.75% 33.56% 28.53% 28.18% 23.67%
difference -11.94% -6.59% -1.63% -1.62% 1.17% -2.90% 1.04%
benefit value $\ell=\infty$ 43.53% 39.76% 34.39% 32.04% 30.22% 25.84% 24.33%
$\ell=2$ 48.63% 45.02% 35.56% 33.39% 28.49% 27.87% 23.42%
difference -5.10% -5.25% -1.16% -1.35% 1.73% -2.03% 0.91%
contribution value $\ell=\infty$ 37.03% 34.58% 32.17% 31.31% 26.83% 25.12% 23.61%
$\ell=2$ 48.74% 47.81% 37.86% 36.47% 30.58% 31.08% 25.45%
difference -$1$3.24% -5.69% -5.16% -3.74% -5.96% -1.84%
Banzhaf* value $\ell=\infty$ 43.90% 39.04% 35.01% 31.64% 29.10% 26.15% 23.91%
$\ell=2$ 47.85% 42.82% 34.13% 32.82% 27.54% 26.71% 22.78%
difference -3.96% -3.78% 0.88% -1.18% 1.56% -0.56% 1.13%
Table 15: Relative ratios for varying country sizes under arbitrary, for $\ell=\infty$, $\ell=2$ and their difference.